OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8 Answer (-2, 0) no solution (3, 4) infinitely many solutions

OpenStudy (anonymous):

do u know matrices?

OpenStudy (anonymous):

@charlieberzak

OpenStudy (anonymous):

ummm no? what is that?

OpenStudy (anonymous):

@charlieberzak Is first equation like this 1/(2x-2y) =4 ?? or is like (1/2x)-2y=4 ??

OpenStudy (anonymous):

i think its 0.5x-2y=4 -x+4y=-8

OpenStudy (anonymous):

from this we can see that line 1 is -2 lots of the second equation. hence the lines are parallel and co-incident. therefore there are infintely many solution.

OpenStudy (anonymous):

-2x0.5=-1 -2x-2=4 -2x4=-8